### v1

@@ 3.3 Part Generation
where $D$ is the pre-trained U-Net model, and $f(\mathbf{p}_{i},\mathbf{a}_{i})$ is the text description of the part $\mathbf{p}_{i}$ constructed using the following process using the part tokens $\mathbf{p}_{i}$ and attributes $\mathbf{a}_{i}$. Intially the the text $f(\mathbf{p}_{i},\mathbf{a}_{i})=\mathbf{p}_{i}$, is set to part token itself. In case the part description (i.e. footnote) is available we set the $f(\mathbf{p}_{i},\mathbf{a}_{i})=\mathbf{a}_{i}^{p}$, further if the style attribute is available we do $f(\mathbf{p}_{i},\mathbf{a}_{i})+=\text{in style of}+\mathbf{a}_{i}^{s}$. In case the color attribute is also specified, the nearest named color $\hat{\mathbf{a}}^{c}_{i}$ (e.g. red) for the specific RBG color $\mathbf{a}_{i}^{c}$ is found, and $f(\mathbf{p}_{i},\mathbf{a}_{i})=\hat{\mathbf{a}}^{c}_{i}+\text{` '}+f(\mathbf{p}_{i},\mathbf{a}_{i})$. The string $f(\mathbf{p}_{i},\mathbf{a}_{i})$ is the text input for the Diffusion to generate the part $\mathbf{p}_{i}$. We use the base prompt $\hat{\mathbf{b}}$ as $f(\mathbf{p}_{i},\mathbf{a}_{i})$ for the background masked region $\mathbf{M}_{b}$. This combination of different diffusion outputs at every time $t$ helps generate a harmonious image after composing the defined parts.
Following Rich-Text [14], we also utilize the gradient guidance [16, 13] by taking the gradient of MSE loss between the estimated original image and the color value $\mathbf{a}_{i}^{c}$ specified by the user. The gradient guidance helps in generating exact color for part $\mathbf{p}_{i}$ which might not be possible with just text guidance [14]. Further, as we use base text framework of Rich-Text, we also can specify font size attribute to $\mathbf{a}_{i}^{f}$ to control the relative importance of each part in the object.
Text-Preservation of Other Parts. As we only intend to modify the object $\mathbf{o}$ from the original prompt, we also use the Self-Injection techniques from Plug and Play [35] to maintain the overall structure of the object from base prompt generation. Further, to make sure that our diffusion trajectory follows the same path as the other object tokens, the background region is also blend with base noise generation.
$x_{t}=\mathbf{M}_{b}\odot x_{t}^{\text{base}}+(1-\mathbf{M}_{b})\odot x_{t}$ (8)
In our case, the attention maps out of the object Mask $\mathcal{M}_{o}$ and all the object parts not assigned to any token comprise the $\mathbf{M}_{b}$, where we start this blend process at $t=T_{blend}$. We find this to be very useful in preserving the structure of the other parts of the image, and just generating the described parts in the localized region. We provide an overview of the full process in Fig. 2.
4 Experimental Analysis
4.1 Part Localization evaluation
@@
Figure 4: Segmentation masks for the parts which are localized by Part Diffusion process, for DeepFashion (above) and CUB200 (below) test dataset images.
We first evaluate the part localization module, which is based on the novel idea of denoising only the object region with the part-based diffusion feature.
Implementation Details. We use the StableDiffusion version 2.1 for our experimentation purpose. We use the DDIM Scheduler [31] with 50 steps to generate results for SD2.1 to evaluate the part localization process. As ground truth is not available for the part masks of the generated images, we use test sets of the commonly used DeepFashion [23] and CUB-200 [38] datasets for evaluation. These datasets are the standard datasets used for the evaluation of the unsupervised part segmentation approaches. The DeepFashion dataset contains images along with their part segmentation masks, divided into 14 categories of labels. The CUB200 dataset contains the key point annotations for the 14 key points specified for the bird categories. We provide further details in Supp. Section.
Baselines and Problem Setting. As we operate in the region of Zero-shot Unsupervised Part Segmentation, there exists no previous works that report results in such a challenging setting. Hence, to facilitate some comparison, we provide results for the unsupervised learning approaches. We provide results for the SCOPS [17], which utilizes the internal features of the VGG model to train a model based on self-supervised loss functions to predict parts robustly across categories. The other stronger self-supervised baseline is Unsup-Parts [11], which uses contrastive loss functions to train a network based on equivariance and other vision properties to cluster the object regions into semantic parts. In addition to this, we also report results for the DFF [10] method on the CUB-200 dataset, as they also operate in the same setting. We want to highlight all these unsupervised approaches to training a neural network on the complete training data to be able to segment parts. In contrast, in comparison, our approach is training free and operates in a zero-shot fashion. Hence, the performance metrics of zero-shot approaches are not directly comparable to unsupervised methods.
Zero Shot Unsupervised Part Segmentation. As there is no benchmark to evaluate the part localization for generated images, we use the existing dataset of DeepFashion and CUB200 to obtain our results. To obtain segmentation for each image, we first invert the image into the diffusion latent space using Null-Text Inversion [26] method (see Sec. 5 for ablation). To obtain approximate text prompts for inversion, we make use of the BLIP-V2 [21] captioner provided in Diffusers. After providing the desired image and prompt, the Null-Text inversion provides us with inverted latent and unconditioned time embeddings to reconstruct the image of the dataset. We then construct a StableDiffusion baseline in which, in addition to the prompts, we also append the list of part tokens $\mathbf{p}$ to the prompt. We then use the segmentation algorithm described above in Sec. 3.2 to extract token maps corresponding to the part tokens $\mathbf{p}_{i}$. We also evaluate a Rich-Text [14] baseline segmentation algorithm for the same. For the proposed part-denoising approach based on parallel diffusion in PartCraft, we first get the inverted latents and embeddings. Then, we generate a base diffusion process to reconstruct the image and use the Part Diffusion to infill the parts of the image. To very fairly compare the StableDiffusion performance with part-denoising, we keep all things the same except the part diffusion process to get masks. Further across these baselines, we use low classifier guidance to ensure proper reconstruction of the dataset image (See Suppl. for details).
@@
Table 1: Comparison to Prior Works for the unsupervised part segmentation task. We follow Unsup-Part [11] for evaluation protocols and baseline results for (K=4) parts.
@@
PartCraft(Ours) 24.7 18.0 48.0 73.4 20.5 9.2 18.5 7.7
Evaluation Protocol. We use the experimental protocol as used by earlier works [11] to facilitate a right point of comparison with earlier works. We want to point out that in our work, the part name (e.g., beak, etc.) is associated with the localized mask, but the unsupervised approaches produce (K=4) parts without any part names. Hence, to make a fair comparison, we create 4 clusters of parts based on their locality, and finally generate segmentation masks with a maximum of 4 clusters. We provide the exact mapping of the part names to cluster labels in the Suppl. Section. We compare the specifically designed metrics [11] of NMI (Normalized Mutual Information) and ARI (Adjusted Rand Index) of the predicted cluster labels with part masks in the case of DeepFashion and key points in case of the CUB dataset. We report the metrics along with their foreground variables (FG-NMI and FG-ARI) for a detailed comparison.
Results. We tabulate the results for all the baselines in Table 1. We find that our approach, PartCraft, significantly improves over the baselines by $\geq$ 10 points in NMI and ARI, both across most metrics in the Unsupervised Zero-Shot setting. Further, we find that our methods, NMI and ARI, are meaningfully high and, in some cases, are near to the Unsupervised learning approaches, too. This demonstrates that non-trivial part masks can be generated by harnessing the power of the pre-trained diffusion models, and they can serve as a good initial prior for learning part segmentation approaches by using them as base pre-trained models. We provide a qualitative results for PartCraft, where we observe that PartCraft is able to associate the right part with the correct label (Fig 4). Meanwhile, in the StableDiffusion baseline, the parts often get assigned to the wrong part tokens, which is the cause of degraded performance (See Supp. Section). We also observe that in the case of our algorithm, if the part gets localized based on the max condition defined in the segmentation algorithm, the part mask found is usually in the right region (See Fig. 4 and Suppl. Sec. for more results). This demonstrates the effectiveness of our part segmentation procedure, which often sides with not localizing incorrect objects rather than performing arbitrary assignment mistakes.
@@
Figure 5: Qualitative Comparison for PartCraft. (above) We provide base prompt ‘a photo of a flamingo’ and part prompt as ‘beak - a pelicans beak’. (below) We generate a photo of a man, with part prompt as ‘black jacket and blue jeans’. We find that other baselines either ignore instruction or change the entire composition of base. On contrary, PartCraft is able to correctly localize part and generate specified details.

### v2

@@ 3.3 Part Generation
where $D$ is the pre-trained U-Net model, and $f(\mathbf{p}_{i},\mathbf{a}_{i})$ is the text description of the part $\mathbf{p}_{i}$ constructed using the following process using the part tokens $\mathbf{p}_{i}$ and attributes $\mathbf{a}_{i}$. Intially the the text $f(\mathbf{p}_{i},\mathbf{a}_{i})=\mathbf{p}_{i}$, is set to part token itself. In case the part description (i.e. footnote) is available we set the $f(\mathbf{p}_{i},\mathbf{a}_{i})=\mathbf{a}_{i}^{p}$, further if the style attribute is available we do $f(\mathbf{p}_{i},\mathbf{a}_{i})=f(\mathbf{p}_{i},\mathbf{a}_{i})+\text{in style of}+\mathbf{a}_{i}^{s}$. In case the color attribute is also specified, the nearest named color $\hat{\mathbf{a}}^{c}_{i}$ (e.g. red) for the specific RBG color $\mathbf{a}_{i}^{c}$ is found, and $f(\mathbf{p}_{i},\mathbf{a}_{i})=\hat{\mathbf{a}}^{c}_{i}+\text{` '}+f(\mathbf{p}_{i},\mathbf{a}_{i})$. The string $f(\mathbf{p}_{i},\mathbf{a}_{i})$ is the text input for the Diffusion to generate the part $\mathbf{p}_{i}$. We use the base prompt $\hat{\mathbf{b}}$ as $f(\mathbf{p}_{i},\mathbf{a}_{i})$ for the background masked region $\mathbf{M}_{b}$. Combining different diffusion outputs at every time $t$ helps generate a harmonious image after blending the defined parts of the object.
@@
Figure 4: Segmentation masks for the parts that are localized by Part Diffusion for DeepFashion (above) and CUB200 (below).
Following Rich-Text [18], we also utilize the gradient guidance [20, 16] by taking the gradient of MSE loss between the estimated original image and the color value $\mathbf{a}_{i}^{c}$ specified by the user. The gradient guidance helps generate the exact RGB color for part $\mathbf{p}_{i}$, which is impossible with just text guidance [18]. Further, as we use the base text framework of Rich-Text, we also can specify font size attribute to $\mathbf{a}_{i}^{f}$ to control the relative importance of each part in the object image.
@@
Table 1: Comparison to Prior Works for the unsupervised part segmentation task. We follow Unsup-Part [12] for evaluation protocols and baseline results for (K=4) parts. We report clustering-based NMI and ARI metrics, which are higher for better segmentation outputs.
@@
PartComposer(Ours) 24.7 18.0 48.0 73.4 20.5 9.2 18.5 7.7
Preservation of Other Parts. As we only intend to modify the object $\mathbf{o}$ from the original prompt, we also use the Self-Injection techniques from Plug and Play [47] to maintain the overall structure of the background from base prompt generation. Further, to ensure that our diffusion trajectory follows the same path as the base, the background region is also blended with base noise generation outputs.
$x_{t}=\mathbf{M}_{b}\odot x_{t}^{\text{base}}+(1-\mathbf{M}_{b})\odot x_{t}$ (5)
In our case, the attention maps out of the object Mask $\mathcal{M}_{o}$ and all the object parts not assigned to any part token comprise the background $\mathbf{M}_{b}$, and we start this blending process at $t=T_{blend}$. We find this to be very useful in preserving the structure of the other parts of the image and just generating the described parts in the localized region. We provide an overview of the complete process in Fig. 2.
4 Experimental Analysis
4.1 Evaluation of Part Localization
We first evaluate the part localization module, which is based on the novel idea of denoising only the object region with the part-based diffusion outputs.
Implementation Details. We use the StableDiffusion (SD) version 2.1 for our experimentation purposes. We use the DDIM Scheduler [43] with 50 steps to generate results for SD2.1 to evaluate the part localization process. As ground truth is not available for the part masks of the generated images, we use test sets of the commonly used DeepFashion [29] and CUB-200 [50] datasets for evaluation. These datasets are the standard datasets used for the evaluation of the unsupervised part segmentation approaches. The DeepFashion dataset contains images along with their part segmentation masks, divided into 14 categories of labels. The CUB200 dataset contains the key point annotations for the 14 key points specified for the bird categories. We provide further details in Suppl. Section.
Baselines and Problem Setting. As we operate in the setting of Zero-shot (i.e. no training) Unsupervised Part Segmentation, there are no previous works that report results in such a challenging setting. Hence, we provide results for the unsupervised learning approaches to facilitate comparison. We provide results for the SCOPS [21], which utilizes the internal features of the VGG model to train a model based on self-supervised loss functions to predict parts robustly across categories. The other stronger self-supervised baseline is Unsup-Parts [12], which uses contrastive loss functions to train a network based on equivariance and other vision properties to cluster the object regions into semantic parts. In addition to this, we also report results for the DFF [11] as they also operate in the same setting. We want to highlight that these unsupervised approaches require either training a neural network or performing clustering on the complete training data to segment parts. In contrast, our approach is training free and operates in a zero-shot fashion. Hence, the performance of zero-shot approaches is not directly comparable to unsupervised methods.
Zero Shot Unsupervised Part Segmentation. As there is no benchmark to evaluate the part localization for generated images, we use the existing dataset of DeepFashion and CUB200 to obtain our results. To obtain segmentation for each image, we first invert the image into the diffusion latent space using Null-Text Inversion [32] method (see Sec. 5 for ablation). We use the BLIP-V2 [25] captioner provided in Diffusers to obtain approximate text prompts for inversion. After providing the desired image and prompt, the Null-Text inversion provides us with inverted latent and unconditioned time embeddings to reconstruct the image of the dataset. We then construct a StableDiffusion baseline in which, in addition to the prompts, we append the list of part tokens $\mathbf{p}$ to the prompt. We then use the segmentation algorithm described above in Sec. 3.2 to extract token maps corresponding to the part tokens $\mathbf{p}_{i}$. We also evaluate a Rich-Text [18] baseline segmentation algorithm for the same. For the proposed part-denoising approach based on parallel diffusion in PartComposer, we first get the inverted latents and embeddings. Then, we generate a base diffusion process to reconstruct the image and use Part Diffusion to fill the parts of the image. To very fairly compare the StableDiffusion performance with part-denoising, we keep all things the same except the part diffusion process to get masks. Further across these baselines, we use low classifier guidance to ensure proper reconstruction of the dataset images (See Suppl. for details).
@@ 4.1 Evaluation of Part Localization
Figure 5: Qualitative Comparison for PartComposer. (above) We provide base prompt ‘a photo of a flamingo’ and part prompt as ‘beak - a pelicans beak’. (below) We generate a photo of a man, with part prompt as ‘black jacket and blue jeans’. We find that other baselines either ignore instruction or change the entire composition. On the contrary, PartComposer can correctly localize parts and generate details.
Evaluation Protocol. We use the standard experimental protocol as used by earlier works [12] to facilitate the right point of comparison. We want to point out that in our work, the part name (e.g., beak, etc.) is associated with the localized mask, but the unsupervised approaches produce (K=4) parts without any part names. Hence, to make a fair comparison, we create 4 clusters of parts based on their locality, and finally generate segmentation masks with a maximum of 4 clusters. We provide the exact mapping of the part names to cluster labels in the Suppl. Section. We compare the specifically designed metrics [12] of NMI (Normalized Mutual Information) and ARI (Adjusted Rand Index) of the predicted cluster labels with part masks in the case of DeepFashion and key points in case of the CUB dataset. We report the metrics along with their foreground variants (FG-NMI and FG-ARI) for a holistic comparison of performance.
Results. We tabulate the results for all the baselines in Table 1. Our approach, PartComposer, significantly improves over the baselines by $\geq$ 5 points in NMI and ARI across most metrics in the Unsupervised Zero-Shot setting. Further, in some cases our methods, NMI and ARI, are near the unsupervised learning approaches. This demonstrates that non-trivial part masks can be generated by harnessing the power of the pre-trained diffusion models, and they can serve as a good initial prior for learning part segmentation approaches by using them as base pre-trained models. We provide qualitative results for PartComposer, where we observe that PartComposer can associate the right part with the correct label (Fig 4). On the contrary, in the StableDiffusion baseline, the parts often get assigned to the wrong part tokens, which is the cause of degraded performance (Suppl. Fig. 16). We also observe that in the case of our algorithm, if the part gets localized based on the max condition defined in the segmentation algorithm, the part mask found is usually in the right region (See Fig. 4 and Table 2). This demonstrates the effectiveness of our part segmentation procedure, which often sides with not localizing objects rather than performing arbitrary assignments.
4.2 PartComposer Image Generation Evaluation
We now evaluate the object composition ability of the proposed approach, PartComposer, compared to the baselines. In this section, we use the base generation model as a StableDiffusion (SD) 1.5 model to do a fair comparison with the Rich-Text baseline. We used the same generation setting of a PNDM [28] scheduler with 41 steps and suggested classifier guidance 8.5. To perform a fair comparison, we keep all the parameters same as that of the Rich-Text [18] baseline.
Baselines. For the task of generating the object image based on part-level details of the specified object, we use the strong baselines of Rich-Text Generation and InstructPix2Pix [8]. The Rich-Text Generation method generates the base image itself, whereas the InstructPix2Pix method requires us to generate the base image. We also evaluate the standard StableDiffusion baseline in which we add all the part details in the text prompt to generate the desired image. We compare all these baselines with our proposed method, PartComposer, while ensuring that the base parameters and seed are the same for the base image generation. A recent work [33] regarding part generation operates in a fully supervised setup using part masks and cannot be used to create novel parts based on text instruction, making it incomparable. We defer specific details for the baselines to the Suppl. Section.
Visual Comparison. We provide a visual comparison for the baselines for the a) bird image distinctive part generation and b) human image part generations. Across both cases (Fig. 5), our proposed method, PartComposer, only modifies the desired beak region and replaces it with the iconic pelican beak. The StableDiffusion baseline modifies the image completely while ignoring instructions (Supp. Fig. 12). In the Rich-Text generation, the full region corresponding to the bird gets modified instead of only the specified prompt in base generation. For the editing-based InstructPix2Pix method, we observe that it modifies the base image at a global level and cannot localize modification to the desired part region (Fig. 5). We further compare and find that recent models like SD3.5, Inpainting [17] and recent SotA editing methods [15, 47, 54, 26] can still not follow exact part instructions as seen in Supp. Fig. 89 and  10. In contrast,observe that PartComposer-generated images are novel aesthetic compositions that follow the part-level instructions.
Quantitative Evaluation. As the part-specific generation is fine-grained, we perform the both automatic and user-study evaluations. We performed a user study by inviting 50 participants (28 responded with a full survey). We evaluate the model by asking questions and evaluating metrics on three orthogonal aspects i.e., a) Localization of part generation in comparison to the base image (through LPIPS [56]), b) Text-Consistency of generated parts by measuring CLIP similarity of the part detailed text prompt with the generated image and c) Aesthetic Quality of the generated image through LAION-5B [42] CLIP Aesthetic scorer. In total, we collect about 3.5k opinions. In Fig. 7, we summarize the results in which we observe that PartComposer is significantly preferred over the other baselines in terms of localized and consistent Part Generation. The automated evaluation results are provided in the table below, which also follow a similar trend as in the user study, demonstrating the effectiveness of PartComposer in a localized part generation while ensuring the aesthetics score is similar to that of the base model.
Table 2: Ablation Analysis of PartComposer for Part Localization, showing performance improvement with each component.
PartComposer (Ablations) Method FG-NMI FG-ARI PartComposer 35.4 11.0 w/o Null-Text Inversion 23.1 5.2 w/o Max Localization 21.3 2.8 w/o Dot Product Localization 23.7 5.0 w/o Independent Text 31.2 8.9 PartComposer (Clustering) PartComposer (K = 9) 35.4 11.0 PartComposer (K = 4) 20.4 2.8 PartComposer (K = 14) 35.2 10.3
Figure 6: PartComposer generalizes over domains as shown by Claude Monet’s painting style (above), where we specify the dress part of women to follow styles like Van Gogh and Ukiyo-e. We perform similar modifications to Woodie’s image.
Figure 7: User Study and Quantitative Results for part-based image generation baselines.
PartComposer StableDiffusion InstructPix2Pix Rich-Text LPIPS 0.168 0.467 0.189 0.243 CLIP 0.201 0.183 0.193 0.187 Aesthetic 5.66 5.68 5.63 5.65
5 Analysis and Discussion
Part Diffusion Based Localization. We ablate the components we have introduced in the PartDiffusion process of PartComposer (Sec. 3.2). We provide an analysis of the effect of using a) null-text inversion, b) max-based localization c) usage of dot product-based protocol in the part assignment, and d) independent text embeddings. We have used a subset of CUB-200 images to perform all evaluations, which is kept fixed across ablations. We tabulate the ablations in Table 2. We observe that all the components introduced in PartComposer contribute significantly to the performance.
Generalization of PartComposer. In Fig. 6, above a painting in Claude Monet style generated by the base StableDiffusion method. We then use PartComposer to specify the dress part of the women’s token, to dark green color in Monet Style, orange color dress in Van Gogh Style (middle), and green color dress in Ukiyo-e style. In Fig. 6, we provide results for the ‘Woody’ character from Toy Story. These results show that PartComposer can generate shirt, pants, and dress color variations, leading to aesthetic image combinations in synthetic and natural (Supp. Fig. 13) domains . This zero-shot generalization across domains demonstrates the creative activities that can be enabled with PartComposer.
6 Conclusion
In this work we introduce PartComposer, a method to generate object images with fine-grained attribute details specified at the part level, using an expressive Rich-Text interface. The PartComposer method introduces a novel part diffusion process, which is responsible for denoising objects using the part features, and then utilizes a region-specific diffusion process to generate part details and compose the final image. PartComposer serves as initial work enabling rich-text-based training-free part-level control for SD models.
Limitations. We find that part generation is bottlenecked with the part understanding. If the part is localized correctly, the text-to-image model can generate specified details in PartComposer. Hence, improving the part-level localization of these models is a good direction for future works.
[…]
Supplementary Material
[…]
Appendix A Additional Qualitative Results
Comparison with Recent Baselines. We compare PartComposer against existing baselines of StableDiffusion 3.5  [17] and Inpainting through Stable Diffusion. For the StableDiffusion (SD) 3.5 model, we use the SD3.5 Large variant for which the web-UI is available online at the following link [^1] https://huggingface.co/spaces/stabilityai/stable-diffusion-3.5-large. We provide the details of the PartComposer by adding those meaningfully into the base prompt. For the swan example in the last row of Fig. 8, we provide the following prompt to the model ‘A photo of a white swan with a peacock crown, 8k, full hd’. We observe in Fig. 8 that although the model tries to adhere to the prompt in ways, it fails to modify the correct object part. Hence, it fails to capture the required details of the part the user expects. For the Inpainting baseline, we use the same StableDiffusion (SD) 2.1 model as a base and use the SAM2 Hiera [39] model for obtaining the segmentation masks. We use the following inpainting repo [^2] https://github.com/Uminosachi/inpaint-anything from GitHub to obtain the results. After obtaining the masks, we provide an inpainting prompt to the SD model to generate the final result. We observe that the SD Inpaint models sometimes, due to incorrect masks for the painting domain, don’t produce results (third row in Fig. 8) and often can’t follow the multiple-part instruction correctly as in Fig. 8 (first row). Further we would like to highlight that inpainting still requires manual ground of masks to parts by user, which makes it a different setting making the quantitative results incomparable. In comparison, PartComposer can produce correct and coherent results by producing the part attributes.
Figure 8: Qualitative Comparison of our proposed method with recent baselines of Stable Diffusion (SD) 3.5 and Stable Diffusion 2.1 Inpainting with SAM2 masks (top right). PartComposer leads to coherent part modification in comparison to advanced baselines.
Comparison with Editing Methods. We provide a comparison of our approach with SotA editing methods: PlugNPlay (PnP) [47], TurboEdit [15], ZONE [26], MagicBrush [54] in Fig. 9. We observe that approaches either do style edits (PnP) or structure edits at full object level (ZONE, MagicBrush). Only PartComposer (ours) can perform the specified localized changes at the semantic part level, demonstrating its novelty and value.
Figure 9: Qualitative Comparison of PartComposer with SotA editing methods applied on the generated base prompt image.
Additional Comparison of PartComposer to Baselines (Fig. 1). In Fig. 11, we demonstrate a comparison of the StableDiffusion, InstructPix2Pix, and Rich-Text on the prompts demonstrated in the teaser figure. For Rich-Text, we add the description in PartText to the part token if it’s present in the base prompt; if not, we add the part description as a footnote in the object token in the base prompt. For InstructPix2Pix, we make instructions regarding part modification one by one on the base generation. We find that the baselines significantly change the entire composition of the object instead of just specified parts.
We also provide additional results for comparison with baselines in Fig. 12, where the Stable Diffusion baseline with part details added often ignores them or generates artifacts. In comparison, PartComposer can produce the desired details for the parts as specified by the user. As in Merida’s example, the Stable Diffusion method doesn’t produce the brown skirt and over uses the green color specified for hair. The Stable Diffusion ignores the desired part detail for the hornbill and cardinal examples. The blue jay example produces an artifact of the dual beak. In comparison, PartComposer generates aesthetic compositions following the part prompt details mentioned by the user.
Appendix B Generalization to SDXL
We implement the PartComposer Method for SDXL, and generate the results for prompts supplied in Fig. 1. Using the SD-XL implementation from Rich-Text Gen [18] as our base code, we implement PartComposer. We present our results in Fig. 10, where we observe that PartComposer can generate high-resolution images with specific attribute details.
Appendix C Attribute Variations
We provide further results for attribute variations for prompts of natural domain in Fig. 13.
Figure 10: High-Resolution PartComposer results using SD-XL as base model, demonstrating its generalization.
Figure 11: Comparison of PartComposer to Other Approaches, for generating using the prompts specified in Fig. 1 of paper. The PartComposer approach can modify base generations in the specified parts and generate aesthetic images compared to the state-of-the-art.
Figure 12: Qualitative Results. We provide a qualitative comparison of additional prompts with Stable Diffusion. We add the part attribute details to the prompt during generation for Stable Diffusion baseline. The PartComposer can generate the image following the specified part attribute details with better coherence and fewer artifacts compared to the baseline.